# Perturbations of Einstein-Maxwell-phantom spacetime: Instabilities of charged Ellis wormholes and quasinormal modes of black holes

### v1

@@ 4.1 Singularity of the gravito-electromagnetic effective potential
introducing the divergent behavior in $V_{11}$. In the EMP spacetime, $\zeta_{\pm}$ mark critical points at which the electromagnetic field transitions from a conventional regime on one side to a phantom-like regime on the other side and can cause the effective potential $V_{11}$ to diverge due to the presence of $Z$ on the denominator. In the Type I black hole scenario, these divergences are strictly confined within the Cauchy horizon ($r_{-}>\zeta_{+}$), a condition equivalent to
@@
Hence, the divergence points of $V_{11}$ are proven to be hidden behind the Cauchy horizon, which indicates the stability of EMP black holes. Nevertheless, as the electric charge increases beyond the extremal value, the black hole undergoes a transition to a wormhole configuration, where singularities of $V_{11}$ become exposed to the outer spacetime, violating cosmic censorship and rendering the configuration physically inadmissible.
@@ 4.2 Finite difference method
$\mathbf{H}\left(t=0,r_{*}\right)=\left(\begin{array}[]{c}\mathrm{e}^{-\frac{\left(r_{*}-a\right)^{2}}{2b^{2}}}\\ \mathrm{e}^{-\frac{\left(r_{*}-a\right)^{2}}{2b^{2}}}\\ 0\end{array}\right),\qquad\mathbf{H}\left(t<0,r_{*}\right)=\mathbf{0},$ (62)
where $a$ and $b$ are the median and width of the initial wave packet, and then calculate the time evolution of the perturbation via the iteration scheme,
@@ 4.3 Divergence of phantom scalar field perturbations
A previous study claimed [78] that EMP wormholes remain stable under massless scalar field perturbations. However, the phantom scalar field, serving as the matter source in the EMP spacetime, carries negative kinetic energy and may therefore induce spacetime instabilities.
Fig. 1 shows how the azimuthal number $l$ affects the effective potential and time domain waveform for phantom scalar field perturbations. For $l=0$, the effective potential contains a negative region, causing the waveform to diverge rapidly after the initial ringing. As $l$ increases, this region disappears and the effective potential develops a double-peak structure, yielding echo signals in the waveform. These echoes follow the patterns remarkably similar to those of a minimally coupled massless scalar field in Ref. [78], and will not be further discussed here. Instead, we focus on the divergence of phantom scalar field perturbations and the singular behavior of $V_{11}$ as mentioned above, both of which indicate the instability of the EMP wormhole. Here we proceed with the following possible explanations:
@@
  • The divergence of $V_{11}$ can be resolved if the Maxwell field vanishes at $r=\zeta_{\pm}$.
@@
As a summary to this section, we have revealed the instability of EMP wormholes from two aspects, where one aspect is the dramatic increment of $l=0$ phantom scalar field perturbations and the other is the divergence of the gravito-electromagnetic effective potential — while the former can be resolved by suitably tuning spacetime parameters, the latter persists. It is the presence of the phantom scalar field and its coupling with the Maxwell field responsible for the instability of EMP wormholes. During the wormhole-black hole transition, the event horizon forms in such a way that it hides these singularities behind the Cauchy horizon. Consequently, although the theory admits wormhole solutions, only black hole configurations with a complete horizon structure are physically viable and stable. A more detailed investigation of the underlying mechanisms and potential remedies remains an open question at present.
5 Quasinormal modes of Type I black holes
@@
$\mathbf{H}\sim\mathrm{e}^{-\mathrm{i}\omega r_{*}},\qquad r_{*}\to-\infty,$ (64)
@@
$\mathbf{H}\sim\mathrm{e}^{\mathrm{i}\omega r_{*}},\qquad r_{*}\to\infty.$ (65)
@@
After assigning the specific electric charge, we can then vary $q$ over the interval $(0,q_{\mathrm{ext}}]$, where $q_{\mathrm{ext}}$ denotes the extremal value,
@@
Figure 2: The effect of the scalar charge $q$ of Type I black holes on the fundamental QNFs of the phantom scalar field perturbation with $\gamma_{1}=1/4$, $\gamma_{2}=2q$, $Q_{\mathrm{e}}/M=2/5$ and $l=2$.
@@
$\tilde{q}=\eta\sin{\vartheta},\qquad\tilde{Q}_{\mathrm{e}}=\eta\cos{\vartheta}.$ ) . (70)
@@
Figs. 3-5 illustrate how the fundamental QNFs depend on the charge composition for the phantom scalar field, electromagnetic field, and gravitational field perturbations of Type I black holes, respectively. The real frequency and damping rate of the perturbed phantom scalar and gravitational fields, see Figs. 3 and 5, decrease monotonically as the mixing angle $\vartheta$ increases, in particular, the amplitudes become smaller for larger values of the generalized specific charge $\eta$. For the perturbed electromagnetic field, see Fig. 4, the real frequency and damping rate show a different tendency: $\omega_{\mathrm{R}}$ increases monotonically as the mixing angle $\vartheta$ increases at a low $\eta$ but decreases monotonically at a high $\eta$; the damping rate rises when the mixing angle is small at a low $\eta$, but it decreases monotonically when the mixing angle increases at a high $\eta$. The values of QNFs are given in Tables 2-4.
@@ 6 Conclusion
The EMP theory can be derived from the EMD theory, where the latter is a four-dimensional low-energy effective theory of string theory. In the EMP spacetime, the phantom scalar field is minimally coupled with Einstein gravity but nonminimally coupled with the Maxwell field. A static and spherically symmetric solution, the so-called charged Ellis wormhole also describes a regular black hole, providing us a frame to study wormholes or black holes under this spacetime background.
We investigate the QNMs of the EMP spacetime by studying the linear perturbation of the metric following Chandrasekhar’s procedure. Because of the spherical symmetry of the EMP spacetime, the radial parts of the perturbed fields can be decomposed from the angular parts, and the latter can be expanded by the spherical harmonics. The perturbation can be classified into the axial (odd-parity) sector and polar (even-parity) sector based on their dependence on the angular coordinate $\varphi$. In the axial sector, the gravitational field perturbation and the electromagnetic field perturbation are coupled, whereas the phantom scalar field perturbation is decoupled. We derive the corresponding three master equations and rewrite the resulting Schrödinger-like equations in a compact matrix form, see Eq. (46). We notice that both the Lagrangian and metric of the EMP theory reduce to general relativity when the phantom scalar hair vanishes and the coupling parameters are chosen properly, so do the master equations and QNFs.
@@
By introducing the generalized specific charge $\eta$ and the mixing angle $\vartheta$, we are able to study the effect of relative contributions of the scalar and electric charges on the QNMs of Type I black holes, see Figs. 3-5 and Tables 2-4. We find that the real frequencies and damping rates behave similarly in the phantom scalar perturbation and the gravitational field perturbation, see Figs. 3 and 5 and Tables 2 and 4 for the details, while they show a different behavior only for a small $\vartheta$ and a low $\eta$ in the electromagnetic field perturbation, see Fig. 4 and Table 3 for the details.
We focus currently on the axial perturbations of the EMP spacetime, where the interesting features on wormholes and black holes have been summarized above. We shall leave the polar perturbations in the EMP spacetime to our future studies and expect additional physical insights.

### v2

@@ 4.1 Singularity of the gravito-electromagnetic effective potential
introducing the divergent behavior in $V_{11}$. In the EMP spacetime, $\zeta_{\pm}$ mark critical points at which the electromagnetic field transitions from a conventional regime on one side to a phantom-like regime on the other side and can cause the effective potential $V_{11}$ to diverge due to the presence of $Z$ on the denominator. Such phantom-like behavior of the electromagnetic field is not intrinsic to the Maxwell 2-form $F_{\mu\nu}$ itself, but is induced entirely by the nonminimal coupling $Z^{-1}$ between the Maxwell field and the phantom scalar field in the Lagrangian. The primary source of the singular behavior of $V_{11}$ is the phantom scalar field: The singular behavior of the gravito-electromagnetic effective potential is a direct consequence of their coupling to the phantom scalar field. If the phantom scalar field were absent, the EMP action would reduce to the Einstein–Maxwell action, and the spacetime would revert to the stable RN solution. In this limit, the singular behavior of the gravito-electromagnetic effective potential would cease to exist.
In the Type I black hole scenario, these divergences are strictly confined within the Cauchy horizon ($r_{-}>\zeta_{+}$), a condition equivalent to
@@
Hence, the divergence points of $V_{11}$ are proven to be hidden behind the Cauchy horizon, which indicates the stability of EMP black holes [^2] Here we address that the linear stability of the exterior region of Type I black holes is studied in analogy with that of RN spacetimes; we note that the interior region of RN black holes, including the extremal case as well, has been shown to be unstable under linear perturbations, see Ref. [102] and the references therein.. Nevertheless, as the electric charge increases beyond the extremal value, the black hole undergoes a transition to a wormhole configuration, where singularities of $V_{11}$ become exposed to the outer spacetime, violating cosmic censorship and rendering the configuration physically inadmissible.
@@ 4.2 Finite difference method
$\mathbf{H}\left(t=0,r_{*}\right)=\left(\begin{array}[]{c}\mathrm{e}^{-\frac{\left(r_{*}-a\right)^{2}}{2b^{2}}}\\ 0\end{array}\right),\qquad\mathbf{H}\left(t<0,r_{*}\right)=\mathbf{0},$ (61)
where $a$ and $b$ are the median and width of the initial wave packet, respectively, and then calculate the time evolution of the perturbation via the iteration scheme,
@@
For the phantom scalar field perturbation, we can simply replace $\mathbf{H}$ and $\mathbf{V}$ in Eq. (58) by $\Psi$ and $V$ in Eq. (53). The initial condition is
$\Psi(t=0,r_{*})=\mathrm{e}^{-\frac{\left(r_{*}-a\right)^{2}}{2b^{2}}},\qquad\Psi(t<0,r_{*})=0.$ (63)
@@ 4.3 Divergence of phantom scalar field perturbations
A previous study reported [80] that EMP wormholes remain stable under massless scalar field perturbations. However, the phantom scalar field, serving as the matter source in the EMP spacetime, carries negative kinetic energy and can therefore induce spacetime instabilities.
Fig. 1 shows how the azimuthal number $l$ affects the effective potential and time domain waveform for phantom scalar field perturbations. For $l=0$, the effective potential contains a negative region, causing the waveform to grow exponentially after the initial ringing. As $l$ increases, this region disappears and the effective potential develops a double-peak structure, yielding echo signals in the waveform. These echoes follow the patterns remarkably similar to those of a minimally coupled massless scalar field in Ref. [80], and will not be further discussed here. Instead, we focus on the divergence of phantom scalar field perturbations and the singular behavior of $V_{11}$ as mentioned above, both of which indicate the instability of the EMP wormhole. Here we proceed with the following possible explanations:
@@
  • The divergence of $V_{11}$ can be resolved if the electromagnetic field perturbation $H_{1}$ vanishes at $r=\zeta_{\pm}$.
@@
The above discussion is based on the linear perturbation theory and does not determine the nonlinear outcome. However, the studies of wormholes supported by phantom scalar fields indicate that the linear instability typically persists nonlinearly, with the spacetime either expanding or collapsing to a black hole depending on the initial perturbation [89, 90, 91]. Whether EMP wormholes follow the same fate or not needs deeper investigations in future.
As a summary to this section, we have revealed two complementary signs of EMP wormhole instability: The exponential growth of $l=0$ phantom scalar field perturbations and a divergence in the gravito-electromagnetic effective potential — while the former can be resolved by suitably tuning spacetime parameters, the latter persists. It is the presence of the phantom scalar field and its coupling with the Maxwell field responsible for the instability of EMP wormholes. During the wormhole–black hole transition, the event horizon forms in such a way that it hides these singularities behind the Cauchy horizon. Consequently, although the theory admits wormhole solutions, only black hole configurations with a complete horizon structure are physically viable and stable. A more detailed investigation of the underlying mechanisms and potential remedies remains an open question for future work.
5 Numerical results
5.1 Quasinormal modes of Type I black holes
@@
$\Psi,\ \mathbf{H}\sim\mathrm{e}^{-\mathrm{i}\omega r_{*}},\qquad r_{*}\to-\infty,$ (64)
@@
$\Psi,\ \mathbf{H}\sim\mathrm{e}^{\mathrm{i}\omega r_{*}},\qquad r_{*}\to\infty,$ (65)
where the tortoise coordinate for Type I black holes is given by
$r_{*}=r+\frac{1}{2\beta}\ln{\frac{\left\lvert r-r_{+}\right\rvert}{2M}}-\frac{\alpha}{2\beta}\ln{\frac{\left\lvert r-r_{-}\right\rvert}{2M}},$ (66)
with
$\beta=\frac{r_{+}-r_{-}}{2\left(r_{+}^{2}+q^{2}\right)},\qquad\alpha=\frac{r_{-}^{2}+q^{2}}{r_{+}^{2}+q^{2}}.$ (67)
The boundary conditions can be approximated by plane waves because the EMP spacetime is asymptotically flat and the perturbation equations take a Schrödinger-like form. The three perturbing fields exhibit similar but not identical asymptotic behavior at the boundaries. To make the distinct asymptotics of each field explicit, we present their Frobenius expansions,
$\psi_{\mathrm{ps}}$ $=\mathrm{e}^{\varrho r}(r-r_{+})^{\tfrac{\varrho}{2\beta}}(r-r_{-})^{-\tfrac{\alpha\varrho}{2\beta}}{\mathrm{e}}^{-2\varrho(r-r_{+})}(r-r_{+}+1)^{-\tfrac{\varrho}{\beta}(1-\alpha)}\sum_{n}a^{\mathrm{ps}}_{n}\Psi^{n},$ (68)
$\psi_{\mathrm{em}}$ $=\mathrm{e}^{\varrho r}(r-r_{+})^{\tfrac{\varrho}{2\beta}}(r-r_{-})^{-\tfrac{\alpha\varrho}{2\beta}}{\mathrm{e}}^{-2\varrho(r-r_{+})}(r-r_{+}+1)^{-\tfrac{\varrho}{\beta}(1-\alpha)}\sum_{n}a^{\mathrm{em}}_{n}H_{1}^{n},$ (69)
$\psi_{\mathrm{g}}$ $=r^{-1}\mathrm{e}^{\varrho r}(r-r_{+})^{\tfrac{\varrho}{2\beta}}(r-r_{-})^{-\tfrac{\alpha\varrho}{2\beta}}{\mathrm{e}}^{-2\varrho(r-r_{+})}(r-r_{+}+1)^{-\tfrac{\varrho}{\beta}(1-\alpha)}\varrho Z^{-1/2}\sum_{n}a^{\mathrm{g}}_{n}H_{2}^{n},$ (70)
where $a_{n}$ denotes the amplitude, and $\psi_{\mathrm{ps}}$, $\psi_{\mathrm{em}}$, and $\psi_{\mathrm{g}}$ correspond to the phantom scalar field, the electromagnetic field and the gravitational field, respectively.
@@
After assigning the specific electric charge, we can then vary $q$ over the interval $(0,q_{\mathrm{ext}}]$. The lower bound $q=0$ follows from the quantum-gravity consistency; the upper bound $q_{\mathrm{ext}}$ denotes the extremal value of black hole configurations, which is fixed in terms of the stability of black hole configurations and the instability of wormhole configurations,
@@
Figure 2: The effect of the scalar charge $q$ of Type I black holes on the QNFs with $\gamma_{1}=1/4$, $\gamma_{2}=2q$, $Q_{\mathrm{e}}/M=2/5$ and $l=2$.
@@
$\tilde{q}=\eta\sin{\vartheta},\qquad\tilde{Q}_{\mathrm{e}}=\eta\cos{\vartheta}.$ (75)
@@
Figs. 3-5 illustrate how the fundamental QNFs depend on the charge composition for the phantom scalar field, electromagnetic field, and gravitational field perturbations of Type I black holes, respectively. The real frequency and damping rate of the perturbed phantom scalar and gravitational fields, see Figs. 3 and 5, decrease monotonically as the mixing angle $\vartheta$ increases, in particular, the decreasing amplitudes increase for larger values of the generalized specific charge $\eta$. For the perturbed electromagnetic field, see Fig. 4, the real frequency and damping rate show a different tendency: $\omega_{\mathrm{R}}$ increases monotonically as the mixing angle $\vartheta$ increases at a low $\eta$ but decreases monotonically at a high $\eta$; the damping rate rises when the mixing angle is small at a low $\eta$, but it decreases monotonically when the mixing angle increases at a high $\eta$. The values of QNFs are given in Tables 2-4. As the parameters approach the transition values from the black-hole side, our numerical survey shows that the quasinormal frequencies vary smoothly and continuously near the transition, suggesting that the spacetime geometry changes rather gently.
@@
5.2 Observability of quasinormal mode deviations
The spectral differences between Type I black holes and their RN counterparts, as shown in Table 1, raise the practical question of whether gravitational wave observations can distinguish the two classes of black holes. A common and heuristic resolvability test for two damped sinusoids is given by the Rayleigh criterion [54],
$(f_{1}-f_{2})\tau>1,$ (76)
where the oscillation frequency $f$ and the damping time $\tau$ are related to the complex QNF by
$f=\frac{\omega_{\mathrm{R}}}{2\pi},\qquad\tau=\frac{1}{\left\lvert\omega_{\mathrm{I}}\right\rvert}$. (77)
The quality factor $\mathscr{Q}$ is defined as
$\mathscr{Q}\equiv\pi f\tau=\frac{\omega_{\mathrm{R}}}{2\left\lvert\omega_{\mathrm{I}}\right\rvert}$, (78)
which characterizes the number of oscillation cycles before the signal decays.
To assess resolvability quantitatively, one must compare the QNM shifts to the measurement uncertainties in frequency $\sigma_{f}$ and damping time $\sigma_{\tau}$. To the leading order and for a large $\mathscr{Q}$, these error scales take the forms [54],
$\sigma_{f}\simeq\frac{1}{\sqrt{2}\pi\tau\iota},\qquad\sigma_{\tau}\simeq\frac{2\tau}{\iota},$ (79)
where $\iota$ denotes the signal-to-noise ratio (SNR) of the ringdown signal. Accordingly, we define the observability measurements for frequency and damping-time deviations as
$\mathcal{O}_{f}\equiv\frac{\left\lvert\Delta f\right\rvert}{\sigma_{f}}=\frac{\left\lvert\Delta f\right\rvert}{f}\frac{f}{\sigma_{f}}=\sqrt{2}\pi\iota\mathscr{Q}\frac{\left\lvert\Delta f\right\rvert}{f},\qquad\mathcal{O}_{\tau}\equiv\frac{\left\lvert\Delta\tau\right\rvert}{\sigma_{\tau}}=\frac{\left\lvert\Delta\tau\right\rvert}{\tau}\frac{\tau}{\sigma_{\tau}}=\frac{\iota}{2}\frac{\left\lvert\Delta\tau\right\rvert}{\tau},$ (80)
where $\Delta f$ and $\Delta\tau$ denote differences between the EMP and RN values. A deviation is considered to be observable when the corresponding ratio exceeds unity, i.e., $\mathcal{O}_{f}>1$ or $\mathcal{O}_{\tau}>1$.
Fig. 6 summarizes our numerical estimates in the following three points:
Figure 6: The observability of deviations in the fundamental gravitational QNM spectra ($l=2,n=0$) between Type I black holes and RN black holes. The blue and red lines represent the observability of frequency $\mathcal{O}_{f}$ and the observability of damping time $\mathcal{O}_{\tau}$ for different values of the scalar charge $q$, with circular, triangular and diamond markers corresponding to SNRs of $\iota=10$, $10^{2}$ and $10^{3}$, respectively. The horizontal line at $\mathcal{O}=1$ indicates the detection threshold.
  • For a modest network SNR $\iota=10$, one representative of most current detections [37], neither the frequency nor the damping-time shift between Type I and RN gravitational QNMs is resolvable, except for the near-extreme cases.
  • For $\iota=10^{2}$ — an SNR that may be achieved for very loud current events [^4] The SNR of the loudest gravitational wave detected to date, GW250114, is 76 [105]. or more commonly by future third-generation ground-based detectors [108, 106, 107] — the frequency deviation becomes resolvable for $q\gtrsim 0.4$, whereas the damping-time deviation is only detectable close to extremality ($q\gtrsim 1.2$).
  • For an optimistic SNR $\iota=10^{3}$, which may be accessible for exceptionally nearby sources with third-generation networks [108, 106, 107], the frequency deviation is distinguishable over most of the $q$ range (except very small $q$), while the damping-time deviation becomes resolvable for $q\gtrsim 0.4$.
These results indicate that the frequency shifts are generally easier to be detected than the changes in damping time, and that the meaningful discrimination of EMP and RN spectra requires either unusually loud events or next-generation detectors. We emphasize several important caveats: The estimates above are based on the Rayleigh heuristic and leading-order error scalings for single-mode, nonspinning ringdowns in an idealized Gaussian noise; they neglect parameter covariances, multi-mode mixing, merger/remnant modelling systematics, detector calibration uncertainties and non-Gaussian noise transients. We also restrict our attention to the fundamental mode and sample only a limited set of SNRs and scalar charges; population statistics, cosmological redshift and full Bayesian parameter estimation have not been included. A more realistic assessment will therefore require injection studies and a Bayesian inference analysis incorporating multi-mode templates and the full detector noise model.
@@ 6 Conclusion
The EMP theory can be derived from the EMD theory, where the latter is a four-dimensional low-energy effective theory of string theory. In the EMP spacetime, the phantom scalar field is minimally coupled with Einstein gravity but nonminimally coupled with the Maxwell field. A static and spherically symmetric solution, the so-called charged EB wormhole also describes a regular black hole, providing us a frame to study wormholes or black holes under this spacetime background.
We investigate the QNMs of the EMP spacetime by studying the linear perturbation of the metric following Chandrasekhar’s procedure. Because of the spherical symmetry of the EMP spacetime, the radial parts of the perturbed fields can be decomposed from the angular parts, and the latter can be expanded by the spherical harmonics. Moreover, the perturbation can be classified into the axial (odd-parity) sector and polar (even-parity) sector based on their dependence on the angular coordinate $\varphi$. Note that the phantom scalar field is of spin-0 (even-parity) and therefore does not enter the axial sector. As a consequence, the axial sector contains only the odd-parity metric perturbations and the odd-parity components of the electromagnetic perturbation, forming a self-contained and coupled gravito-electromagnetic system. In the polar sector, the additional physical degrees of freedom emerge, leading to a more complex and physically significant gravito-electromagnetic-phantom scalar coupled system. In order to study the stability of the spacetime under phantom scalar field perturbations only, where the phantom scalar field perturbations are a primary source of potential instability, we adopt a controlled approximation in the polar sector. We next derive the corresponding master equations and rewrite the resulting Schrödinger-like equations in a compact matrix form, see Eq. (42). We notice that both the Lagrangian and metric of the EMP theory reduce to general relativity when the phantom scalar hair vanishes and the coupling parameters are chosen properly, so do the master equations and QNFs.
@@
By introducing the generalized specific charge $\eta$ and the mixing angle $\vartheta$, we are able to study the effect of relative contributions of the scalar and electric charges on the QNMs of Type I black holes, see Figs. 35 and Tables 24. We find that the real frequencies and damping rates behave similarly in the phantom scalar perturbation and the gravitational field perturbation, see Figs. 3 and 5 and Tables 2 and 4 for the details, while they show a different behavior only for a small $\vartheta$ and a low $\eta$ in the electromagnetic field perturbation, see Fig. 4 and Table 3 for the details.
Our investigation into the observability of QNM deviations yields a preliminary but encouraging assessment. Although the spectral shifts between Type I and RN black holes are theoretically significant, they are small: The frequency shifts are more readily measurable than changes in the damping time (see Fig. 6). Given the current gravitational-wave catalog, the robust constraints on the EMP model appear unlikely; nevertheless, future high-precision ringdown spectroscopy — particularly with third-generation detectors — could provide a direct empirical test for phantom scalar hairs, transforming these spectral signatures from theoretical predictions into viable observational phenomena.
We focus currently on the linear perturbations of the EMP spacetime, where the interesting features on wormholes and black holes have been summarized above. We shall leave the polar perturbations in the EMP spacetime and a comprehensive observability study to future work and expect new physical insights.